OpenStudy (anonymous):

Write the quadratic variation equation for the relationship: y varies directly with x2 and y = 36 when x = 3. A. y = ¼x^2 B. y = 12x C. y = 4x^2 D. y = x^2 + 25

OpenStudy (anonymous):

I think its B

OpenStudy (jack1):

sub 3 into each of the ones with x^2 in it, what comes up with 36?

OpenStudy (tanner23456):

the standard formula for directly related is: y = kx in this case they want y = kx^2 so if x = 3 and y = 36 then 36 = k (3)^2 now solve for k divide the 9 k = 36/9 k = 4 therefore y = 4x^2